# Transferable Model-agnostic Vision-Language Model Adaptation for Efficient Weak-to-Strong Generalization

### v2

@@ 1 Introduction
The rapid evolution of vision-language models (VLMs) (Jia et al. 2021; Radford et al. 2021) has driven significant advancements in computer vision. These models are typically built with two modality-specific encoders, one for image and one for text, and are trained on large datasets of image-text pairs using contrastive loss. Here, modality-specific encoders learn to map each modality input into a joint embedding space, enabling zero-shot capabilities across a wide range of visual recognition tasks. The broad and generalized knowledge embedded in VLMs enables efficient adaptation on downstream tasks with few labeled data as explored in (Yang et al. 2025; Wang et al. 2024; Zhou et al. 2022b, a; Khattak et al. 2023; khattak et al. 2023; Li et al. 2024). However, as VLMs are scaled (Chen et al. 2024; Cherti et al. 2023) or equipped with more complex architectures (Yu et al. 2022), fine-tuning them requires massive memory and computational costs.
Figure 1: Comparison of adaptation knowledge transfer methods. Performance is averaged over 11 visual recognition tasks in a base-to-novel setting. TransMiter outperforms other adaptation transfer approaches, including Prompt Transfer (Su et al. 2022) and EFT (Mitchell et al. 2024), while maintaining inference speed nearly identical to a zero-shot model (gray), which serves as the upper-bound. With a small amount of labeled data, TransMiter+ surpasses its supervised counterpart (Khattak et al. 2023), e.g., Fine-tuning (red).
In this context, an important research question arises; “How can the knowledge gained from adapting smaller, weaker models, i.e., adaptation knowledge, be effectively leveraged to improve larger and more advanced models?” This challenge is well known as weak-to-strong generalization (Burns et al. 2024). One common approach (Su et al. 2022; Han et al. 2024; Liu et al. 2024a; Mitchell et al. 2024) involves directly reusing a subset of parameters from a weaker model to a stronger one. However, due to semantic discrepancy or dimensional mismatch between different models, additional training is often necessary (Su et al. 2022; Han et al. 2024) to align their internal representations properly. Some works (Lu et al. 2023; Liu et al. 2024a; Mitchell et al. 2024; Lu et al. 2024) mitigate this issue by enforcing semantic consistency (e.g., model prediction), reducing the need for re-training but at the cost of slower inference due to multiple model usage. Alternatively, knowledge distillation (Yuan et al. 2020; Nasser, Gupte, and Sethi 2024; Burns et al. 2024; Li et al. 2024) has been demonstrated as an effective way for smaller models to guide larger ones. Still, this approach requires expensive retraining whenever a new model is introduced, limiting its practicality for rapidly evolving models.
@@ 2 Related Works
Vision-Language Model Adaptation. Vision-language models (VLMs), such as CLIP (Radford et al. 2021), have demonstrated strong generalization abilities across various visual recognition tasks. Leveraging this strength, numerous studies have investigated fine-tuning VLMs on visual recognition tasks, such as prompt learning (Yang et al. 2025; Wang et al. 2024; Zhou et al. 2022b, a; Zhu et al. 2023; Khattak et al. 2023; khattak et al. 2023; Li et al. 2024), low-rank adaptation (Zanella and Ayed 2024), linear probing (Ouali et al. 2023), and regularization (Zhu et al. 2023; Khattak et al. 2023). However, existing VLM adaptation methods often neglect the models’ rapid evolution in terms of size and complexity, resulting in the inconvenience of retraining (Su et al. 2022) whenever a new model comes. To address this limitation, we focus on developing a transferable adaptation module for VLMs, enabling the efficient transfer of adaptation knowledge acquired during fine-tuning across different VLMs, regardless of their size or architectural design.
Weak-to-Strong Generalization. In an era of rapidly evolving VLMs, “weak-to-strong generalization”, where stronger models leverage the knowledge of weaker models, is emerging as a new research challenge for enabling efficient model adaptation. Previous methods try to transfer either a portion of a model’s internal parameters (Su et al. 2022; Chijiwa 2024; Han et al. 2024) or the entire model itself (Lu et al. 2023; Liu et al. 2024a; Mitchell et al. 2024) to other models. However, differences in architecture often hinder parameter reuse, and even when feasible, aligning representation spaces requires costly retraining. While full-model transfer alleviates the problem using predicted logits, it requires running multiple models, leading to significant inference cost. Some other studies (Yuan et al. 2020; Nasser, Gupte, and Sethi 2024; Burns et al. 2024; Li et al. 2024) have shown that weaker models can effectively supervise stronger ones, a concept known as reverse knowledge distillation (Nasser, Gupte, and Sethi 2024). While these methods incur no additional inference cost, they still require expensive retraining whenever a new model is introduced, making it impractical for frequent model updates. In our work, we design an adaptation module for high cross-model transferability and fast inference speed. Additionally, we employ “knowledge extraction and transfer” approach, where adaptation knowledge is extracted only “once” from a weaker model and efficiently transferred to a stronger model.
@@ 4 Method
4.1 Transferable Model Agnostic Adapter
@@
Previous adaptation techniques (Zhou et al. 2022b; khattak et al. 2023; Zhou et al. 2022a; Zhu et al. 2023; Khattak et al. 2023; Li et al. 2024; Ouali et al. 2023) rely on internal model parameters that differ across VLMs, making direct transfer across different VLMs challenging. To mitigate this problem, our adapter uses the predicted logits from the VLM as input. The logits exhibit two key properties: (1) a fixed dimensionality that aligns with the number of classes, and (2) semantic consistency for each logit element. These properties facilitate effective communication (Moschella et al. 2023) between different models, enabling seamless reuse of our adapter. Moreover, our adapter does not require access to the parameters of VLM, thereby achieving high efficiency.
@@
A key consideration is adapting to the task while preserving the generalized knowledge in the original model. Since TransMiter employs simple 2D matrices to facilitate transitions between the logit and latent space, we enforce an inverse relationship between the projection and reconstruction matrices. If the transformation function $f$ operates as an identity function (e.g., , $f(h)=h$), the output remains identical to the original logits, making an inverse relationship between the matrices reasonable. Here, we adopt orthogonality for the two matrices, i.e., $W_{\text{in}}W_{\text{in}}^{\intercal}=W_{\text{out}}^{\intercal}W_{\text{out}}=I$, and set the two matrices in a transposed relationship, i.e., $W_{\text{in}}=W_{\text{out}}^{\intercal}$. These properties provide the projection and reconstruction matrices with an implicit regularization effect (Liu et al. 2021; Bansal, Chen, and Wang 2018; Qiu et al. 2023) and establish an inverse relationship, i.e., $W_{\text{in}}W_{\text{out}}=I$, between them. For simplicity, we will denote $W_{\text{out}}$ as $W_{\text{s}}$, representing the transition matrix for the weak VLM. We can rewrite equation 5 in a simplified form as follows:
@@ 4.2 Forward-only Adapter Transfer
Table 1: Performance on base-to-base adaptation transfer. We combine source and target models among RN50, ViT-B/16, and ViT-L-14. Gray-colored rows represent the performance of the target model when fine-tuned, serving as the upper bound.
@@
where $\mathcal{O}=\{W\in\mathbb{R}^{D\times D},W^{\intercal}W=I_{D}\}$ denotes the orthogonality constraint.
The solution can be written as:
@@
Here, the projection matrix for strong VLM is updated to $W_{\text{s}}^{\intercal}\hat{W}$. Likewise, the reconstruction matrix for strong VLM is updated to $\hat{W}^{\intercal}W_{\text{s}}$. By doing so, the inverse relationship between the projection and reconstruction matrices and their orthogonality are preserved. This process ensures that strong VLM retains its pre-trained knowledge while absorbing adaptation knowledge from weak VLM. Here, we can compute $\hat{W}^{\intercal}W_{\text{s}}$ as a single transition matrix $W_{\text{t}}$ for strong VLM. The final outputs can be written as:
@@ 5.1 Experimental setup
Settings. During the adaptation knowledge transfer, we need a fine-tuned weak VLM. Based on how this model was fine-tuned, we set up two main evaluation setups: (1) Base-to-base adaptation transfer, where the weak model is trained on all the task classes, and (2) Base-to-novel adaptation transfer, where the weak model is fine-tuned on a subset of classes (i.e., seen classes), and tested on both seen and the remaining classes (i.e., unseen classes). Through these configurations, we aim to evaluate how effectively the methods transfer task-specific adaptation knowledge from weaker to stronger models while preserving its generalized knowledge. Details of the experimental settings are in the Supplement.
Datasets. We evaluate our method on 11 visual recognition datasets for both base-to-base and base-to-novel adaptation transfer settings. For fine-tuning the models, we follow CoOp (Zhou et al. 2022b) and CoCoOp (Zhou et al. 2022a), using a 16-shot setup. All unlabeled image samples (Li et al. 2024) are used for training and transfer in TransMiter. See Supplement for more details.
Baselines. We adopt CLIP (Radford et al. 2021) for vision-language models, varying with architecture and size, e.g., ResNet-50 (RN50), ViT-B/16, and ViT-L/14. We assume that the target model has stronger generalization capabilities than the source model in terms of its size and architecture. For fine-tuning the source model, we apply CoOp (Zhou et al. 2022b) and PromptSRC (Khattak et al. 2023). In each setting, we evaluate three methods within our framework: (1) Source Fine-tuning $\theta_{\text{ft-s}}$, where the source model is fine-tuned; (2) Target Zero-shot $\theta_{\text{pt-t}}$, the zero-shot performance of the pre-trained target model; and (3) Target TransMiter $\theta_{\text{t}}^{*}$, where TransMiter is applied to the pre-trained target model. As described in Section 3, our objective is to improve the strong model beyond the performance of the fine-tuned weak model (Source Fine-tuning) and pre-trained strong model (Target Zero-shot). Additionally, we report the performance of Target Fine-tuning $\theta_{\text{ft-t}}$, where the strong model is fine-tuned. It serves as the ‘upper bound’.
To compare with transferable adaptation methods, we employ the following baselines: EFT (Mitchell et al. 2024; Liu et al. 2024a) and Prompt Transfer (Su et al. 2022). Prompt Transfer transfers learned soft prompts from weaker to stronger models. This approach requires resource-intensive training to align the representation spaces between weak and strong models. To achieve this, we adopt a naive knowledge distillation method (Li et al. 2024; Hinton, Vinyals, and Dean 2015). In contrast, EFT operates in a training-free manner but relies on multiple weak models to assist the strong model during inference, leading to high inference costs. Additional details are provided in the Supplement.
@@ 5.2 Experimental Results
Base-to-base adaptation transfer. Table 1 reports the performance of our method and baselines across 11 visual recognition datasets under the base-to-base adaptation transfer setting. To evaluate the inference efficiency of each method, we also reported the additional GFLOPs required compared to the base model.
@@
In addition to the unsupervised adaptation transfer setting, we further fine-tune TransMiter with labeled data (TransMiter+) to fully assess its capability, and compare it against recent state-of-the-art supervised fine-tuning approaches in Table 3. We integrate additional adaptation strategies, such as Maple (khattak et al. 2023), HPT (Wang et al. 2024), and LwEIB (Yang et al. 2025), into TransMiter for a weaker model (e.g., ViT-B/16) adaptation, which serves as a ‘fine-tuned weak VLM’ during the adaptation knowledge extraction process. Even without the adaptation knowledge extraction phase (row 2), TransMiter+ leads to a substantial improvement over the zero-shot baseline. Moreover, since TransMiter’s knowledge extraction and transfer provide a strong initialization, fine-tuning TransMiter with supervision after the adaptation knowledge transfer consistently yields improved performance across various strategies. Importantly, combining any fine-tuning strategy with TransMiter not only surpasses the performance of its fine-tuning counterpart but also lowers inference cost, demonstrating both high effectiveness and efficiency.
@@ 5.3 Ablation study
Table 4: Ablation study of TransMiter. PA, ACE, and BS refer to “Prediction-based Adapter”, “Auxiliary Class Expansion”, and “Basis Change”, respectively.
In this section, we evaluate the effectiveness of TransMiter through ablations. Unless otherwise stated, we use ViT-B/16 and ViT-L/14 as the source and target VLMs, respectively, and PromptSRC (Khattak et al. 2023) for fine-tuning.
@@
Model $M$ DTD ($\Delta$) UCF ($\Delta$) Cars ($\Delta$) SUN ($\Delta$)
47 100 196 397
Source $N_{\text{task}}$ 69.82, 0.00 83.84, 0.00 79.62, 0.00 76.28, 0.00
256 71.77, +1.95 84.30, +0.46 79.96, +0.34 -
512 72.32, +2.50 84.32, +0.48 79.87, +0.25 76.07, -0.21
1024 72.54, +2.72 84.37, +0.53 80.52, +0.90 76.46, +0.18
Target $N_{\text{task}}$ 64.18, 0.00 81.96, 0.00 81.38, 0.00 76.52, 0.00
256 71.91, +7.73 84.25, +2.29 81.47, +0.09 -
512 73.64, +9.46 85.15, +3.19 83.03, +1.65 76.79, +0.27
1024 74.21, +10.03 85.52, +3.56 83.11, +1.73 77.14, +0.62
Effect of anchor quantity. In Table 5, we examine the effects of anchor quantity $M$ on both source and target models within a base-to-base setting. We observe that increasing $M$ enhances the performance of both the source and target models. Notably, the performance gain from additional anchors for the target model is relatively higher across all reported datasets than for the source model. Further, the performance increases with more anchors. These results underscore two main implications: (1) incorporating auxiliary classes enhances representational capacity, leading to more effective training of TransMiter and (2) they play a crucial role in achieving high transferability.
@@
Efficiency analysis. In Table 6, we evaluate the computational costs of TransMiter on StanfordCars dataset. Among unsupervised transfer methods, TransMiter achieves strong performance and fast inference, while its backpropagation-free design significantly reduces transfer time and memory. With a small amount of labeled data, TransMiter+ surpasses the performance and inference speed of the fine-tuning strategy (PromptSRC) used in its development. This improvement takes less than a minute of extra training, as it updates only the adapter on top of the VLM, whereas PromptSRC requires almost full-model gradient computation, highlighting TransMiter’s superior efficiency.

### v3

@@ 1 Introduction
The rapid evolution of vision-language models (VLMs) (Jia et al. 2021; Radford et al. 2021) has driven significant advancements in computer vision. These models are typically built with two modality-specific encoders, one for image and one for text, and are trained on large datasets of image-text pairs using contrastive loss. Here, modality-specific encoders learn to map each modality input into a joint embedding space, enabling zero-shot capabilities across a wide range of visual recognition tasks. The broad and generalized knowledge embedded in VLMs enables efficient adaptation on downstream tasks with few labeled data, as explored in (Yang et al. 2025b; Wang et al. 2024; Zhou et al. 2022b, a; Khattak et al. 2023b, a; Li et al. 2024). However, as VLMs are scaled (Chen et al. 2024; Cherti et al. 2023) or equipped with more complex architectures (Yu et al. 2022), fine-tuning them requires massive memory and computational costs.
Figure 1: Comparison of adaptation knowledge transfer methods. Performance is averaged over 11 visual recognition tasks in a base-to-novel setting. TransMiter (ours) outperforms other adaptation transfer approaches, including Prompt Transfer (Su et al. 2022) and EFT (Mitchell et al. 2024), while maintaining inference speed nearly identical to a zero-shot model (gray), which serves as the upper-bound. With a small amount of labeled data, TransMiter+ (ours+) surpasses its supervised counterpart (Khattak et al. 2023b), e.g., Fine-tuning (red).
In this context, an important research question arises; “How can the knowledge gained from adapting smaller, weaker models, i.e., adaptation knowledge, be effectively leveraged to improve larger and more advanced models?” One common approach (Su et al. 2022; Han et al. 2024; Liu et al. 2024a; Mitchell et al. 2024) involves directly reusing a subset of parameters from a weaker model to a stronger one. However, due to semantic discrepancy or dimensional mismatch between different models, additional training is often necessary (Su et al. 2022; Han et al. 2024) to align their internal representations properly. Some works (Lu et al. 2023; Liu et al. 2024a; Mitchell et al. 2024; Lu et al. 2024) mitigate this issue by enforcing semantic consistency (e.g., model prediction), reducing the need for re-training but at the cost of slower inference due to multiple model usage. Alternatively, knowledge distillation (Yuan et al. 2020; Nasser et al. 2024; Burns et al. 2024; Li et al. 2024) has been demonstrated as an effective way for smaller models to guide larger ones. Still, this approach requires expensive retraining whenever a new model is introduced, limiting its practicality for rapidly evolving models.
@@ 2 Related Works
Vision-Language Model Adaptation. Vision-language models (VLMs), such as CLIP (Radford et al. 2021), have demonstrated strong generalization abilities across various visual recognition tasks. Leveraging this strength, numerous studies have investigated fine-tuning VLMs on visual recognition tasks, such as prompt learning (Yang et al. 2025b; Wang et al. 2024; Zhou et al. 2022b, a; Zhu et al. 2023; Lee et al. 2023; Cho et al. 2023; Khattak et al. 2023b, a; Li et al. 2024; Yang et al. 2025a), low-rank adaptation (Zanella and Ayed 2024), linear probing (Ouali et al. 2023), and regularization (Park et al. 2024; Zhu et al. 2023; Khattak et al. 2023b). However, existing VLM adaptation methods often neglect the models’ rapid evolution in terms of size and complexity, resulting in the inconvenience of retraining (Su et al. 2022) whenever a new model comes. To address this limitation, we focus on developing a transferable adaptation module for VLMs, enabling the efficient transfer of adaptation knowledge acquired during fine-tuning across different VLMs, regardless of their size or architectural design.
Weak-to-Strong Generalization. In an era of rapidly evolving VLMs, “weak-to-strong generalization”, where stronger models leverage the knowledge of weaker models, is emerging as a new research challenge for enabling efficient model adaptation. Previous methods try to transfer either a portion of a model’s internal parameters (Su et al. 2022; Chijiwa 2024; Han et al. 2024) or the entire model itself (Lu et al. 2023; Liu et al. 2024a; Mitchell et al. 2024) to other models. However, differences in architecture often hinder parameter reuse, and even when feasible, aligning representation spaces requires costly retraining. While full-model transfer alleviates the problem using predicted logits, it requires running multiple models, leading to significant inference cost. Some other studies (Yuan et al. 2020; Nasser et al. 2024; Burns et al. 2024; Li et al. 2024) have shown that weaker models can effectively supervise stronger ones, a concept known as reverse knowledge distillation (Nasser et al. 2024). While these methods incur no additional inference cost, they still require expensive retraining whenever a new model is introduced, making it impractical for frequent model updates. In our work, we design an adaptation module for high cross-model transferability and fast inference speed. Additionally, we employ “knowledge extraction and transfer” approach, where adaptation knowledge is extracted only “once” from a weaker model and efficiently transferred to a stronger model.
@@ 4 Method
4.1 Transferable Model-Agnostic Adapter
@@
Previous adaptation techniques (Zhou et al. 2022b; Khattak et al. 2023a; Zhou et al. 2022a; Zhu et al. 2023; Khattak et al. 2023b; Li et al. 2024; Ouali et al. 2023) rely on internal model parameters that differ across VLMs, making direct transfer across different VLMs challenging. To mitigate this problem, our adapter uses the predicted logits from the VLM as input. The logits exhibit two key properties: (1) a fixed dimensionality that aligns with the number of classes, and (2) semantic consistency for each logit element. These properties facilitate effective communication (Moschella et al. 2023) between different models, enabling seamless reuse of our adapter. Moreover, our adapter does not require access to the parameters of VLM, thereby achieving high efficiency.
@@
A key consideration is adapting to the task while preserving the generalized knowledge in the original model. Since TransMiter employs simple 2D matrices to facilitate transitions between the logit and latent space, we enforce an inverse relationship between the projection and reconstruction matrices. If the transformation function $f$ operates as an identity function (e.g., , $f(h)=h$), the output remains identical to the original logits, making an inverse relationship between the matrices reasonable. Here, we adopt orthogonality for the two matrices, i.e., $W_{\text{in}}W_{\text{in}}^{\intercal}=W_{\text{out}}^{\intercal}W_{\text{out}}=I$, and set the two matrices in a transposed relationship, i.e., $W_{\text{in}}=W_{\text{out}}^{\intercal}$. These properties provide the projection and reconstruction matrices with an implicit regularization effect (Liu et al. 2021; Bansal et al. 2018; Qiu et al. 2023) and establish an inverse relationship, i.e., $W_{\text{in}}W_{\text{out}}=I$, between them. For simplicity, we will denote $W_{\text{out}}$ as $W_{\text{s}}$, representing the transition matrix for the weak VLM. We can rewrite (5) in a simplified form as follows:
@@ 4.2 Forward-only Adapter Transfer
Table 1: Performance on base-to-base adaptation transfer. We combine source and target models among RN50, ViT-B/16, and ViT-L/14. Gray-colored rows represent the performance of the target model when fine-tuned, serving as the upper bound.
@@
where $\mathcal{O}=\{W\in\mathbb{R}^{D\times D},W^{\intercal}W=I_{D}\}$ denotes the orthogonality constraint. The mapping matrix $\hat{W}$ can be derived as the solution to the Orthogonal Procrustes problem (Schönemann 1966) via singular value decomposition, making backpropagation unnecessary. The solution can be written as:
@@
Here, the projection matrix for strong VLM is updated to $W_{\text{s}}^{\intercal}\hat{W}$. Likewise, the reconstruction matrix for strong VLM is updated to $\hat{W}^{\intercal}W_{\text{s}}$. By doing so, the inverse relationship between the projection and reconstruction matrices and their orthogonality are preserved. This process ensures that strong VLM retains its pre-trained knowledge while absorbing adaptation knowledge from weak VLM. The final outputs of the enhanced strong VLM can be written as:
@@ 5.1 Experimental setup
Settings. During the adaptation knowledge transfer, we need a fine-tuned weak VLM. Based on how this model was fine-tuned, we set up two main evaluation setups: (1) Base-to-base adaptation transfer, where the weak model is trained on all the task classes, and (2) Base-to-novel adaptation transfer, where the weak model is fine-tuned on a subset of classes (i.e., seen classes), and tested on both seen and the remaining classes (i.e., unseen classes). Through these configurations, we aim to evaluate how effectively the methods transfer task-specific adaptation knowledge from weaker to stronger models while preserving their generalized knowledge. We evaluate our method on 11 visual recognition datasets for both settings. See Supplement for more details.
Baselines. We adopt CLIP (Radford et al. 2021) for VLMs, varying with architecture and size, e.g., ResNet-50 (RN50), ViT-B/16, and ViT-L/14. We assume that the target model has stronger generalization capabilities than the source model in terms of its size and architecture. For fine-tuning the source model, we apply CoOp (Zhou et al. 2022b) and PromptSRC (Khattak et al. 2023b). In each setting, we evaluate three methods within our framework: (1) Source Fine-tuning $\theta_{\text{ft-s}}$, where the source model is fine-tuned; (2) Target Zero-shot $\theta_{\text{pt-t}}$, the zero-shot performance of the pre-trained target model; and (3) Target TransMiter $\theta_{\text{t}}^{*}$, where TransMiter is applied to the pre-trained target model. As described in Section 3, our objective is to improve the strong model beyond the performance of the fine-tuned weak model (Source Fine-tuning) and pre-trained strong model (Target Zero-shot). Additionally, we report the performance of Target Fine-tuning $\theta_{\text{ft-t}}$, where the strong model is fine-tuned. It serves as the ‘upper bound’.
To compare with transferable adaptation methods, we employ the following baselines: EFT (Mitchell et al. 2024; Liu et al. 2024a) and Prompt Transfer (Su et al. 2022). Prompt Transfer transfers learned soft prompts from weaker to stronger models. This approach requires resource-intensive training to align the representation spaces between weak and strong models. To achieve this, we adopt a naive knowledge distillation method (Li et al. 2024; Hinton et al. 2015). In contrast, EFT operates in a training-free manner but relies on multiple weak models to assist the strong model during inference, leading to high inference costs. Additional details are provided in the Supplement.
@@ 5.2 Experimental Results
Base-to-base adaptation transfer. Table 1 reports the performance of our method and baselines across 11 visual recognition datasets under the base-to-base adaptation transfer setting. To evaluate the inference efficiency of each method, we also reported the additionally introduced GFLOPs compared to the base model.
@@
In addition to the unsupervised adaptation transfer setting, we further fine-tune TransMiter with labeled data (TransMiter+) to fully assess its capability, and compare it against recent state-of-the-art supervised fine-tuning approaches in Table 3. We integrate additional adaptation strategies, such as MaPLe (Khattak et al. 2023a), HPT (Wang et al. 2024), and LwEIB (Yang et al. 2025b), into TransMiter for a weaker model (e.g., ViT-B/16) adaptation, which serves as a ‘fine-tuned weak VLM’ during the adaptation knowledge extraction process. Even without the adaptation knowledge extraction phase (row 2), TransMiter+ leads to a substantial improvement over the zero-shot baseline. Moreover, since TransMiter’s knowledge extraction and transfer provide a strong initialization, fine-tuning TransMiter with supervision after the adaptation knowledge transfer consistently yields improved performance across various strategies. Importantly, combining any fine-tuning strategy with TransMiter not only surpasses the performance of its fine-tuning counterpart but also lowers inference cost, demonstrating both high effectiveness and efficiency.
@@ 5.3 Ablation study
Table 4: Ablation study of TransMiter. PA, ACE, and BC refer to “Prediction-based Adapter”, “Auxiliary Class Expansion”, and “Basis Change”, respectively.
In this section, we evaluate the effectiveness of TransMiter through ablations. Unless otherwise stated, we use ViT-B/16 and ViT-L/14 as the source and target VLMs, respectively, and PromptSRC (Khattak et al. 2023b) for fine-tuning.
@@
Efficiency analysis. In Table 5, we evaluate the computational costs of TransMiter on StanfordCars dataset. Among unsupervised transfer methods, TransMiter achieves strong performance and fast inference, while its backpropagation-free design significantly reduces transfer time and memory. With a few labeled data, TransMiter+ surpasses the performance and inference speed of the fine-tuning strategy (PromptSRC) used in its development. This improvement takes less than a minute of extra training, as it updates only the adapter on top of the VLM, whereas PromptSRC requires almost full-model gradient computation, highlighting TransMiter’s superior efficiency.